# LayoutCoT: Unleashing the Deep Reasoning Potential of Large Language Models for Layout Generation

### v1

@@ 3.2 Coarse Layout Generation
After retrieving the top-$K$ examples via RAG (Section 3.1), we convert each selected layout ^$\hat{\mathcal{L}}_{k}$ into a pre-defined HTML format. Let $\mathcal{R}_{K}(\mathcal{L})=\{\hat{\mathcal{L}}_{1},\dots,\hat{\mathcal{L}}_{K}\}$ be the retrieved set; we encode each ^$\hat{\mathcal{L}}_{k}$ as an HTML snippet and combine all of them with the given constraint $c$. Formally, we provide the large language model (LLM) with:
$\Bigl{\{}\mathrm{HTML}(\hat{\mathcal{L}}_{1}),\,\dots,\,\mathrm{HTML}(\hat{\mathcal{L}}_{K}),\,c\Bigr{\}},$
where $\mathrm{HTML}(\hat{\mathcal{L}}_{k})$ denotes the HTML representation of layout ^$\hat{\mathcal{L}}_{k}$, and $c$ specifies user-defined constraints.
We then prompt the LLM to generate $n$ candidate layouts
$\mathcal{L}^{1}_{HTML}=\text{LLM}\Bigl{(}\{\mathrm{HTML}(\hat{\mathcal{L}}_{1}),\,\dots,\,\mathrm{HTML}(\hat{\mathcal{L}}_{K}),\,c\}\Bigr{)}.$ (4)
By repeating the above generation process $n$ times, we obtain
$\Bigl{\{}\mathcal{L}^{1}_{HTML},\,\dots,\,\mathcal{L}^{n}_{HTML}\Bigr{\}},$
each in HTML format. To select the most promising one from these $n$ candidates, we employ a ranker $\tau(\cdot)$ following LayoutPrompter [28]. Specifically, we compute a layout quality score $\tau\bigl{(}\mathcal{L}^{j}_{HTML}\bigr{)}$ for each generated layout $\mathcal{L}^{j}_{HTML}$ and pick the highest-scoring layout as our coarse result:
[…]
This coarse layout, $\mathcal{L}^{\mathrm{coarse}}_{HTML}$, preserves the key structural constraints specified by the user while serving as the initial configuration for subsequent refinements.
Figure 3: Details of the CoT Module. We illustrate the overall conversational logic of the multi-stage CoT. Depending on the type of task, there are slight variations in the details to adapt to specific requirements.
@@ 3.3 Chain of Thought for Layout Refinement
Once we have obtained the coarse layout $\mathcal{L}^{\mathrm{coarse}}_{HTML}$ under constraint $c$, we employ a multi-stage Chain-of-Thought (CoT) strategy to iteratively refine the layout. Concretely, we break the entire layout generation task (still under the user-defined constraint $c$) into $i$ simpler refinement subproblems, denoted by $\{Q_{L}^{1},Q_{L}^{2},\dots,Q_{L}^{i}\}$. Each $Q_{L}^{t}$ ($1\leq t\leq i$) focuses on a specific aspect of improving the layout (e.g., rearranging bounding boxes, reducing overlaps, or adjusting spacing among elements).
Iterative Refinement Procedure. At the first refinement step ($t=1$), we feed the following information to the large language model (LLM):
$\Bigl{\{}\mathrm{HTML}(\hat{\mathcal{L}}_{1}),\dots,\mathrm{HTML}(\hat{\mathcal{L}}_{K}),\;\mathcal{L}^{\mathrm{coarse}}_{HTML},\;c,\;Q_{L}^{1}\Bigr{\}},$
where $\{\mathrm{HTML}(\hat{\mathcal{L}}_{1}),\dots,\mathrm{HTML}(\hat{\mathcal{L}}_{K})\}$ are the top-$K$ prompting examples (cf. Section 3.1), $\mathcal{L}^{\mathrm{coarse}}_{HTML}$ is the initial coarse layout, $c$ is the user-defined constraint, and $Q_{L}^{1}$ is the first refinement subproblem. The LLM then outputs:
$\mathcal{L}_{HTML}^{\mathrm{refine}_{1}}=\text{LLM}\Bigl{(}\mathrm{HTML}(\hat{\mathcal{L}}_{1}),\dots,\mathcal{L}^{\mathrm{coarse}}_{HTML},c,\;Q_{L}^{1}\Bigr{)}.$ (6)
At the $t$-th refinement step ($1<t\leq i$), the input is updated with the newly refined layout from the previous step, $\mathcal{L}_{HTML}^{\mathrm{refine}_{t-1}}$, together with the same top-$K$ examples, constraint $c$, and the next subproblem $Q_{L}^{t}$:
$\Bigl{\{}\mathrm{HTML}(\hat{\mathcal{L}}_{1}),\dots,\mathrm{HTML}(\hat{\mathcal{L}}_{K}),\;\mathcal{L}_{HTML}^{\mathrm{refine}_{t-1}},\;c,\;Q_{L}^{t}\Bigr{\}},$
yielding:
$\mathcal{L}_{HTML}^{\mathrm{refine}_{t}}=\text{LLM}\Bigl{(}\mathrm{HTML}(\hat{\mathcal{L}}_{1}),\dots,\mathcal{L}_{HTML}^{\mathrm{refine}_{t-1}},c,\;Q_{L}^{t}\Bigr{)}.$ (7)
After completing all $i$ subproblems, the final refined layout is:
$\mathcal{L}_{HTML}^{\mathrm{refine}_{i}}=\text{LLM}\Bigl{(}\mathrm{HTML}(\hat{\mathcal{L}}_{1}),\dots,\mathcal{L}_{HTML}^{\mathrm{refine}_{i-1}},c,\;Q_{L}^{i}\Bigr{)}.$
This multi-stage, question-and-answer refinement process allows the LLM to tackle each localized challenge in turn, resulting in a more coherent and aesthetically pleasing final layout.
Table 1: Dataset statistics.
Dataset Tasks Training Set Test Set Element Types
PKU content-aware layout generation 9,974 905 3
CGL content-aware layout generation 38,510 1,647 5
RICO constraint-explicit layout generation 31,694 3,729 25
PubLayNet constraint-explicit layout generation 311,397 10,998 5
WebUI text-to-layout 3,835 487 10
Table 2: Content-Aware layout generation task results, all obtained under unconstrained generation conditions. LayoutCoT denotes the results when using ’DeepSeek-R1’ as the model with the CoT module removed, while both LayoutPrompter and LayoutCoT use ’gpt-4’ as the model.
CGL PKU
Method Training-free Content Graphic Content Graphic
[…]
CGL-GAN [47] 0.489 0.268 0.147 0.0420 0.275 0.244 0.269 0.876 - 0.219 0.175 0.226 0.0062 0.575 0.259 0.0605 0.707 -
DS-GAN [14] 0.451 0.224 0.193 0.0485 0.370 0.301 0.075 0.893 - 0.209 0.217 0.229 0.0046 0.585 0.424 0.026 0.903 -
RALF [13] 0.336 0.197 0.247 0.0023 0.943 0.884 0.027 0.995 0.874 0.171 0.150 0.246 0.0007 0.893 0.840 0.018 0.998 0.981
LayoutPrompter [28] 0.415 0.184 0.227 0.0091 0.338 0.317 0.0049 0.996 0.812 0.251 0.171 0.237 0.0021 0.824 0.809 0.0005 0.997 0.865
LayoutCoT 0.267 0.182 0.225 0.0018 0.515 0.501 0.0023 0.994 0.803 0.214 0.180 0.264 0.0016 0.925 0.849 0.0004 1.000 0.806
LayoutCoT (ours) 0.170 0.101 0.260 0.0005 0.958 0.958 0.0018 0.998 0.939 0.207 0.074 0.289 0.0002 0.980 0.946 0.0013 1.000 1.000
@@ 4.1 Experimental Setup
We conduct detailed experiments on three different types of tasks: content-aware, constraint-explicit, and text-to-layout, using five distinct datasets: PKU [14], CGL [46], RICO [29], Publaynet [45], and WebUI [40]. Since no official test set splits are provided for the PKU and CGL datasets, we used 9,974 labeled images in the PKU dataset as the training set and the database for the Layout-aware RAG module, with 905 unlabeled images serving as the test set. In the CGL dataset, we took 38,510 images as the training set and the database for the Layout-aware RAG module, and selected an additional 1,647 images (with no overlap with the training set) as the test set. For other datasets, we adopt the official data splits, as detailed in Table .1
@@ 4.2 Implement Details
In our experiments, we provide details on the choice of the $k$-value for the top-$k$ retrieval in the Layout-aware RAG. When generating the Coarse Layout, we follow the approach of LayoutPrompter [28] and set $k$=10. For the CoT stage, we set $k$=4. We also adopt LayoutPrompter’s setting of $n$=10 for the number of generation times by LLMs when selecting the Coarse Layout. Due to the ’Text-Davinci-003’ model used by LayoutPrompter is no longer available, we replace the LLM with ’gpt-4’ for comparison. Furthermore, when conducting the LayoutCoT experiments, we also use ’gpt-4’ as the LLM. In addition, we will provide the performance results of ’DeepSeek-R1’ after removing the CoT module.
@@ 4.3 Evaluation Metrics
For the constraint-explicit layout generation and text-to-layout tasks, we utilize four standard metrics for evaluation: Alignment measures how well the elements within the layout align with each other. Overlap calculates the area of overlap between any two elements in the layout. Maximum IoU, mIoU assesses the maximum intersection over union between the generated layout and the actual layout, providing a measure of layout accuracy. Frechet Inception Distance (FID) quantifies the similarity between the distribution of generated layouts and the distribution of real layouts, using features extracted from a deep neural network trained on images. For the content-aware layout generation task, we reference the eight metrics used in [14] along with a newly designed metric $R_{e}$ that we introduce. The metric $R_{e}$ is used to assess the reasonableness of each element’s size within the layout. Let
$r_{i}\;=\;\frac{S^{i}}{S^{i}_{\text{train}}},\qquad d_{i}\;=\;\bigl{|}\ln r_{i}\bigr{|},\qquad\tau\;=\;\ln(1.1),$
where $S^{i}$ is the predicted average area of $label_{i}$ and $S^{i}_{\text{train}}$ is the corresponding training‑set reference. The constant $\tau$ encodes a ±$\pm 10\%$ tolerance band.
$\operatorname{score}_{i}\;=\;\exp\!\bigl{(}-\max(0,\,d_{i}-\tau)\bigr{)}.$ (8)
Eqa.(8) yields $\operatorname{score}_{i}=1$ whenever $0.9\leq r_{i}\leq 1.1$ and decays exponentially (symmetrically for oversized and undersized elements) once the deviation exceeds 10%.
For $N$ labels, the overall area‑error metric is
$R_{e}^{coarse}\;=\;\frac{1}{N}\sum_{i=1}^{N}\operatorname{score}_{i}.$ (9)
A more extreme‑error‑sensitive alternative is
$R_{e}\;=\;\exp\!\Bigl{(}-\sqrt{\frac{1}{N}\sum_{i=1}^{N}\bigl{(}\max(0,\,d_{i}-\tau)\bigr{)}^{2}}\Bigr{)}.$ (10)
Both $R_{e}^{coarse}$ and $R_{e}$ lie in $(0,1]$.
@@
4.4 Results on Content-Aware Layout Generation
The quantitative experimental results are shown in Table 2. As seen from the table, our method surpasses all the evaluation metrics used in content-aware layout generation tasks compared to both the training-based baseline methods and another training-free method, LayoutPrompter [28]. Notably, our proposed CoT module enables ”GPT-4” to even outperform the deep reasoning large language model ”DeepSeek-R1” in this layout generation task. Moreover, our performance on the Overlap metric in the PKU dataset was suboptimal, prompting further visual analysis. As illustrated in Figure. 4, the very low scores on the Overlap metric for both LayoutPrompter and LayoutCoT result from numerous bad cases. In these cases, layout bounding boxes are extremely small and densely clustered in the upper-left corner, leading to low Overlap values but highly unaesthetic and impractical layout outcomes. Hence, we introduce the $R_{e}$ metric in Sec. 4.3 to measure the reasonableness of layout sizes, on which we have performed exceptionally well. Qualitative comparative cases are further provided in Figure. 4.
Table 3: Quantitative comparison with other methods on constraint-explicit layout generation tasks. LayoutCoT denotes the results when using ’DeepSeek-R1’ as the model with the CoT module removed.
RICO PubLayNet
Tasks Methods mIoU $\uparrow$ FID $\downarrow$ Align $\downarrow$ Overlap $\downarrow$ mIoU $\uparrow$ FID $\downarrow$ Align $\downarrow$ Overlap $\downarrow$
Gen-T BLT [22] 0.216 25.633 0.150 0.983 0.140 38.684 0.036 0.196
LayoutFormer++ [19] 0.432 1.096 0.230 0.530 0.348 8.411 0.020 0.008
LayoutPrompter [28] 0.604 2.134 0.0018 0.0285 0.619 1.648 0.0009 0.0037
LayoutCoT 0.718 4.970 0.0015 0.0208 0.795 2.608 0.0001 0.0027
LayoutCoT (Ours) 0.719 2.189 0.0008 0.0220 0.772 1.227 0.0001 0.0024
Gen-TS BLT [22] 0.604 0.951 0.181 0.660 0.428 7.914 0.021 0.419
LayoutFormer++ [19] 0.620 0.757 0.202 0.542 0.471 0.720 0.024 0.037
LayoutPrompter [28] 0.707 3.072 0.0020 0.0505 0.727 1.863 0.0007 0.076
LayoutCoT 0.782 1.252 0.0005 0.0331 0.844 1.701 0.0001 0.034
LayoutCoT (Ours) 0.795 3.506 0.0004 0.0305 0.847 1.430 0.0003 0.004
Gen-R CLG-LO [21] 0.286 8.898 0.311 0.615 0.277 19.738 0.123 0.200
LayoutFormer++ [19] 0.424 5.972 0.332 0.537 0.353 4.954 0.025 0.076
LayoutPrompter [28] 0.583 2.249 0.0007 0.0436 0.627 2.224 0.0008 0.011
LayoutCoT 0.724 2.399 0.0006 0.0301 0.779 1.853 0.0002 0.004
LayoutCoT (Ours) 0.725 3.569 0.0009 0.0145 0.769 1.408 0.0005 0.002
Completion LayoutTransformer [11] 0.363 6.679 0.194 0.478 0.077 14.769 0.019 0.0013
LayoutFormer++ [19] 0.732 4.574 0.077 0.487 0.471 10.251 0.020 0.0022
LayoutPrompter [28] 0.541 2.656 0.0010 0.0405 0.754 1.282 0.0009 0.0008
LayoutCoT 0.682 2.236 0.008 0.0309 0.843 0.840 0.0007 0.0003
LayoutCoT (Ours) 0.716 2.112 0.0008 0.0117 0.770 0.531 4e-5 0.0006
Refinement RUITE [33] 0.811 0.107 0.133 0.483 0.781 0.061 0.029 0.020
LayoutFormer++ [19] 0.816 0.032 0.123 0.489 0.785 0.086 0.024 0.006
LayoutPrompter [28] 0.874 0.225 0.0011 0.1095 0.756 0.143 0.003 0.0006
LayoutCoT 0.869 0.876 0.0006 0.0714 0.801 0.265 0.001 0.0002
LayoutCoT (Ours) 0.890 0.353 0.0003 0.0520 0.839 0.141 0.0004 0.0003

### v3

@@ 3.2 Coarse Layout Generation
After retrieving the top-$K$ examples via RAG (Section 3.1), we convert each layout $\hat{\mathcal{L}}_{k}\!\in\!\mathcal{R}_{K}(\mathcal{L})$ into a pre‑defined HTML snippet, the format definition of the HTML snippets is consistent with that of LayoutPrompter. Let the unified prompt set be
$\mathcal{S}\;=\;\bigl\{\mathrm{HTML}(\hat{\mathcal{L}}_{1}),\,\dots,\,\mathrm{HTML}(\hat{\mathcal{L}}_{K}),\;c\bigr\}$ (4)
where $c$ contains user‑defined constraints. Feeding $\mathcal{S}$ to the LLM yields $n$ candidate layouts in parallel:
$\bigl\{\mathcal{L}^{j}_{\mathrm{HTML}}\bigr\}_{j=1}^{n}=\text{LLM}(\mathcal{S})$ (5)
We score each candidate with a ranker $\tau(\cdot)$ Lin et al. (2023b) and select the highest‑scoring one as the coarse result:
$\mathcal{L}^{\mathrm{coarse}}_{\mathrm{HTML}}\;=\;\arg\max_{\,1\leq j\leq n}\;\tau\bigl(\mathcal{L}^{j}_{\mathrm{HTML}}\bigr)$ (6)
This coarse layout preserves the user’s key structural constraints and serves as the starting point for refinement.
@@ 3.3 Chain of Thought for Layout Refinement
Given $\mathcal{L}^{\mathrm{coarse}}_{\mathrm{HTML}}$ and the same constraint $c$, we apply a multi‑stage Chain‑of‑Thought strategy to iteratively improve the design. Let the task be decomposed into $i$ refinement sub‑problems $\{Q_{L}^{1},\dots,Q_{L}^{i}\}$. Denote $\mathcal{L}^{\mathrm{refine}_{0}}_{\mathrm{HTML}}=\mathcal{L}^{\mathrm{coarse}}_{\mathrm{HTML}}$. At stage $t$ ($1\!\leq\!t\!\leq\!3$) we query the LLM with the augmented prompt $\mathcal{S}\cup\{\mathcal{L}^{\mathrm{refine}_{t-1}}_{\mathrm{HTML}},\,Q_{L}^{t}\}$:
$\mathcal{L}^{\mathrm{refine}_{t}}_{\mathrm{HTML}}=\text{LLM}\Bigl(\mathcal{S}\cup\bigl\{\mathcal{L}^{\mathrm{refine}_{t-1}}_{\mathrm{HTML}},\,Q_{L}^{t}\bigr\}\Bigr),\,t=1,2,3$
The final refined layout $\mathcal{L}^{\mathrm{refine}_{3}}_{\mathrm{HTML}}$ is thus obtained after $3$ sequential reasoning steps, yielding a design that is both semantically coherent and visually pleasing.
@@ 4.1 Experimental Setup
We conduct detailed experiments on three different types of tasks: content-aware, constraint-explicit, and text-to-layout, using five distinct datasets: PKU Hsu et al. (2023), CGL Zhou et al. (2022a), RICO Liu et al. (2018), Publaynet Zhong et al. (2019), and WebUI Wu et al. (2023). Since no official test set splits are provided for the PKU and CGL datasets, we used 9,974 labeled images in the PKU dataset as the training set and the database for the Layout-aware RAG module, with 905 unlabeled images serving as the test set. In the CGL dataset, we took 38,510 images as the training set and the database for the Layout-aware RAG module, and selected an additional 1,647 images (with no overlap with the training set) as the test set. For other datasets, we adopt the official data splits, as detailed in Table .1
@@ 4.2 Implement Details
In our experiments, we provide details on the choice of the $k$-value for the top-$k$ retrieval in the Layout-aware RAG. When generating the Coarse Layout, we follow the approach of LayoutPrompter Lin et al. (2023b) and set kk=10. For the CoT stage, we set kk=4. We also adopt LayoutPrompter’s setting of nn=10 for the number of generation times by LLMs when selecting the Coarse Layout. Since the ’Text-Davinci-003’ model used by LayoutPrompter is no longer available, we replaced the LLM with ’gpt-4’ in our comparative experiments. Additionally, on the PKU dataset, we conducted extra experiments using Gemini-2.5-pro, denoted as LayoutPrompter. In the LayoutCoT experiments, we used ’gpt-4’ as the LLM. To further validate the effectiveness of our CoT module, we provided LayoutCoT results for each experiment; LayoutCoT refers to the variant where the LLM is replaced by the reasoning model ’DeepSeek-R1’ and the CoT module is removed (only LayoutAware RAG and Coarse Layout Generation are used). For the PKU dataset, we also conducted experiments using Gemini-2.5-pro, denoted as LayoutCoT.
Figure 4: Qualitative Results for the Content-aware Layout Generation Task. LayoutPrompter and LayoutCoT tend to generate dense, small boxes in the upper-left corner, which is unreasonable. In contrast, LayoutCoT effectively corrects this error and achieves satisfactory results.
@@ 4.3 Evaluation Metrics
For the constraint-explicit layout generation and text-to-layout tasks, we utilize four standard metrics for evaluation: Alignment measures how well the elements within the layout align with each other. Overlap calculates the area of overlap between any two elements in the layout. Maximum IoU, mIoU assesses the maximum intersection over union between the generated layout and the actual layout, providing a measure of layout accuracy. Frechet Inception Distance (FID) quantifies the similarity between the distribution of generated layouts and the distribution of real layouts, using features extracted from a deep neural network trained on images. For the content-aware layout generation task, we reference the eight metrics used in Hsu et al. (2023) along with a newly designed metric $R_{e}$ that we introduce. The metric $R_{e}$ is used to assess the reasonableness of each element’s size within the layout. Let
$r_{i}=\frac{S^{i}}{S^{i}_{\text{train}}},\qquad d_{i}=\bigl|\ln r_{i}\bigr|,\qquad\tau=\ln(1.1)$
where $S^{i}$ is the predicted mean area of label $i$ and $S^{i}_{\text{train}}$ is its training-set reference. With a $\pm 10\%$ tolerance band $\tau$, we define
$\operatorname{score}_{i}=\exp\,\!\bigl(-\max(0,\,d_{i}-\tau)\bigr)$ (8)
Thus $\operatorname{score}_{i}=1$ when $0.9\leq r_{i}\leq 1.1$ and decays exponentially once the deviation exceeds $10\%$.
Aggregating over all $N$ labels yields
$R_{e}=\exp\Bigl(-\sqrt{1/N\sum\nolimits_{i=1}^{N}\bigl(\max(0,\,d_{i}-\tau)\bigr)^{2}}\Bigr)$ (9)
where $R_{e}\in(0,1]$; higher values indicate more reasonable element sizes.
4.4 Results on Content-Aware Layout Generation
The quantitative experimental results for both PKU and CGL datasets are summarized in Table 2.. As seen from the table, our method surpasses all the evaluation metrics used in content-aware layout generation tasks compared to both the training-based baseline methods and another training-free method, LayoutPrompter Lin et al. (2023b). Notably, our proposed CoT module enables "GPT-4" to even outperform the deep reasoning large language model "DeepSeek-R1" in this layout generation task. Moreover, our performance on the Overlap metric in the PKU dataset was suboptimal, prompting further visual analysis. As illustrated in Figure. 4, the very low scores on the Overlap metric for both LayoutPrompter and LayoutCoT result from numerous bad cases. In these cases, layout bounding boxes are extremely small and densely clustered in the upper-left corner, leading to low Overlap values but highly unaesthetic and impractical layout outcomes. Hence, we introduce the $R_{e}$ metric in Sec. 4.3 to measure the reasonableness of layout sizes, on which we have performed exceptionally well. Qualitative comparative cases are further provided in Figure. 4.
Table 3: Quantitative comparison with other methods on constraint-explicit layout generation tasks.
RICO PubLayNet
@@ 4.4 Results on Content-Aware Layout Generation
Gen-T BLT Kong et al. (2022) 0.216 25.63 0.150 0.983 0.140 38.68 0.036 0.196
LayoutFormer++ Jiang et al. (2022) 0.432 1.096 0.230 0.530 0.348 8.411 0.020 0.008
LayoutPrompter Lin et al. (2023b) 0.604 2.134 0.002 0.029 0.619 1.648 0.001 0.004
LayoutCoT 0.718 4.970 0.002 0.021 0.795 2.608 0.000 0.003
LayoutCoT (Ours) 0.719 2.189 0.001 0.022 0.772 1.227 0.000 0.002
Gen-TS BLT Kong et al. (2022) 0.604 0.951 0.181 0.660 0.428 7.914 0.021 0.419
LayoutFormer++ Jiang et al. (2022) 0.620 0.757 0.202 0.542 0.471 0.720 0.024 0.037
LayoutPrompter Lin et al. (2023b) 0.707 3.072 0.002 0.051 0.727 1.863 0.001 0.076
LayoutCoT 0.782 1.252 0.001 0.033 0.844 1.701 0.000 0.034
LayoutCoT (Ours) 0.795 3.506 0.000 0.031 0.847 1.430 0.000 0.004
Gen-R CLG-LO Kikuchi et al. (2021) 0.286 8.898 0.311 0.615 0.277 19.73 0.123 0.200
LayoutFormer++ Jiang et al. (2022) 0.424 5.972 0.332 0.537 0.353 4.954 0.025 0.076
LayoutPrompter Lin et al. (2023b) 0.583 2.249 0.001 0.044 0.627 2.224 0.001 0.011
LayoutCoT 0.724 2.399 0.001 0.030 0.779 1.853 0.000 0.004
LayoutCoT (Ours) 0.725 3.569 0.001 0.015 0.769 1.408 0.001 0.002
Comp. LayoutTrans Gupta et al. (2021) 0.363 6.679 0.194 0.478 0.077 14.76 0.019 0.001
LayoutFormer++ Jiang et al. (2022) 0.732 4.574 0.077 0.487 0.471 10.25 0.020 0.002
LayoutPrompter Lin et al. (2023b) 0.541 2.656 0.001 0.041 0.754 1.282 0.001 0.001
LayoutCoT 0.682 2.236 0.008 0.031 0.843 0.840 0.001 0.000
LayoutCoT (Ours) 0.716 2.112 0.001 0.012 0.770 0.531 0.000 0.001
Refine RUITE Rahman et al. (2021) 0.811 0.107 0.133 0.483 0.781 0.061 0.029 0.020
LayoutFormer++ Jiang et al. (2022) 0.816 0.032 0.123 0.489 0.785 0.086 0.024 0.006
LayoutPrompter Lin et al. (2023b) 0.874 0.225 0.001 0.110 0.756 0.143 0.003 0.001
LayoutCoT 0.869 0.876 0.001 0.071 0.801 0.265 0.001 0.000
LayoutCoT (Ours) 0.890 0.353 0.000 0.052 0.839 0.141 0.000 0.000